Circular No. 7389

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000au IN MCG +8-15-14 T. Puckett and A. Langoussis, Mountain Town, GA, report the dis-

covery of an apparent supernova (mag 17.9) on an unfiltered CCD frame (limiting mag 19.3) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Mar. 28.16 UT. SN 2000au is located at α = 7^{h}53^{m}08^{s}.40, δ = +50^{o}07′35′′.3 (equinox 2000.0), which is 5′′.1 west and 14′′.5 south of the center of MCG +8-15-14. SN 2000au was also present on an unfiltered CCD frame taken on Mar. 29.05, but it does not appear on an unfiltered image taken on Feb. 29.15, neither does the object appear on Palomar Sky Survey images taken on 1998 Mar. 24 (limiting mag ∼ 21.0) and 1954 Dec. 30 (limiting mag ∼ 20.0).

XTE J1118+480 R. Remillard, E. Morgan, D. Smith, and E. Smith report for the RXTE

ASM team at the Massachusetts Institute of Technology and at Goddard Space Flight Center, NASA: “The RXTE All-Sky Monitor has detected x-ray emission from a new source, XTE J1118+480. The ASM position is centered at α = 11^{h}18^{m}17^{s}, δ = +48^{o}03′.0 (equinox 2000.0; 90-percent-confidence radius 6′). The average x-ray intensity (2–12 keV) was 39(8) mCrab on Mar. 29. Retrospective ASM analysis indicates that the x-ray flux has been slowly rising since Mar. 5. Another modest outburst had previously occurred during Jan. 2–29, peaking at 37(3) mCrab on Jan. 6. The SIMBAD database lists only an unremarkable guide star (V ∼ 11) near the x-ray position. The source was confirmed in an 800-s RXTE pointed observation beginning on Mar. 29.968 UT. Rapid flares (e.g., 10 s) were seen, reaching a factor of 5 above the baseline flux near 25 mCrab and suggesting a Galactic x-ray source. No pulsations were detected. The x-ray spectrum does not vary appreciably during intense flares. Preliminary spectral analysis indicates a power law visible to 30 keV. The photon index is roughly 1.8, which is similar to Cyg X-1 in its hard state. The high galactic latitude (+62^{o}) and x-ray properties present a puzzling combination, and observations at other wavelengths are urgently needed.”

SUPERNOVA 2000D IN UGC 1767 Unfiltered CCD magnitudes: Feb. 5.406 UT, 17.8: (S. Yoshida and K.

Kadota, Ageo, Saitama, Japan); Mar. 13.76, 17.5 ± 0.2 (K. Sárneczky, K. Sziládi, and L. Kiss, Piszkéstető station, Konkoly Observatory).

2000 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green